# Improved High-probability Convergence Guarantees of Decentralized SGD

### v1

@@ 5 Conclusion
In this paper we revisit convergence in HP of a variant of DSGD under sub-Gaussian noise. We show that DSGD is guaranteed to converge in the HP sense under the same conditions on the cost as in the MSE sense, achieving order-optimal rates for both non-convex and strongly convex costs. Moreover, our results show DSGD achieves linear speed-up in the number of users in both cases. Compared to [57], wherein the authors study HP convergence of DSGD, we relax strong conditions like uniformly bounded gradients or asymptotically vanishing noise, while improving the rates by showing linear speed-up in the number of users. Compared to works studying HP convergence of centralized SGD, e.g., [31, 32, 33, 40], we provide tight control on the MGF of the consensus gap for both non-convex and strongly convex costs, in Lemmas 3 and 4, as well as a more fine-grained bound on the MGF of an “almost decreasing” process, in Lemma 5. Future work includes extending our results to costs satisfying the PL condition, incorporating the GT mechanism to remove bounded heterogeneity condition and considering noise with heavier tails, like sub-Weibull.
@@ Appendix A Introduction
The Appendix contains results omitted from the main body. Section B collects some important facts used in our proofs, Section C provides some technical results, Section D defines some notions used in the analysis, Sections E and F provide proofs for non-convex and strongly convex costs, respectively, while Section G contains further discussion on dimension dependence in our bounds.
@@ Appendix B Useful Inequalities
Next, we state a useful result from [25].
Proposition 6 ([25], Lemma 21).
@@
The following result states some important consequences of conditions (A2) and (A3), see, e.g., [76, 74].
@@
Finally, we complete this section with the following important consequences of (A6), see, e.g., [74].
@@
Let (A6) hold. Then, the following are true, for all $i\in[n]$ and $x,y\in\mathbb{R}^{d}$.
@@ Appendix C Technical Results
    To prove the first property, we follow steps similar to those in [39, Lemma 1]. Let $y_{i}=\frac{z^{t}_{i}}{\sigma_{i}}$ and note that $\mathbb{E}\left[\exp\left(\|y_{i}\|^{2}\right)\>|\>\mathcal{F}_{t}\right]\leq\exp(1)$, from (A3). Assume first that $v\in\mathbb{R}^{d}$ is such that $\|v\|\leq\frac{4}{3}$. Using the inequality $\exp(a)\leq a+\exp(\nicefrac{{9a^{2}}}{{16}})$, which holds for any $a\in\mathbb{R}$, we then have
@@
    Combining the previous result with Lemma 1 from [75], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
@@
    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., [77, Proposition 2.5.2], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
@@ Appendix D Analysis Setup
We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [13, 18, 25]. Let ${\mathbf{x}}^{t}\coloneq col(x_{1}^{t},\ldots,x_{n}^{t})\in\mathbb{R}^{nd}$ denote the column vector stacking users’ local models. Using this notation, we can then represent the update rule (12) compactly as
@@
where $\mathbf{W}=W\otimes I_{d}\in\mathbb{R}^{nd\times nd}$, $\otimes$ denotes the Kronecker product and $I_{d}\in\mathbb{R}^{d\times d}$ denotes the $d$-dimensional identity matrix, while ${\mathbf{g}}^{t}\coloneqq col(g^{t}_{1},\ldots,g^{t}_{n})$. Next, define the matrix $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$, where $\mathbf{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones. The matrix $J$ represent the “ideal” consensus matrix, where all users can communicate with each other. [^11] This is equivalent to the client-server setup in terms of the update rule, as the server averages models from all clients in each iteration. The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [18, 25] and references therein.
@@ Appendix E Proofs for Non-convex Costs
The proof follows similar steps as in [25, Lemma 3]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
@@
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$, where $C>0$ is a problem related constant satisfying
@@
$\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+L\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}[\Delta_{x}+n\lambda^{2}(A^{2}+\sigma^{2})]}{(1-\lambda^{2})T}\Bigg).$
@@
    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{dT}}$, then with probability at least $1-\delta$
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\sigma L\sqrt{d}}{\sqrt{nT}}+\frac{48n\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{dT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{dT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma}{\sqrt{nT}}+\frac{1}{CT}$. Therefore, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we finally have
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+120L\big)}{\sqrt{nT}}$
    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{48n\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
@@
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{T+1}}{{\delta}})\big)}{\sqrt{n(T+1)}}+\frac{\sigma\sqrt{d}\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
    as the last term, which does not attain a linear speed-up, is no longer of higher order. We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [11, 25].
@@ Appendix F Proofs for Strongly Convex Costs
In this section we prove Lemma 4 and Theorem 2. To do so, we follow a similar strategy to the one in, e.g., [20, 25], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense. However, to establish guarantees in the HP sense, we instead work with the MGF of the iterates and have the following important result.
@@
$\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+d\sigma^{2}\kappa/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+\kappa^{2})}}{{(1-\lambda)}})}{(1-\lambda)n(T+t_{0})^{2}}$
$+\frac{\Delta_{f}}{n(T+t_{0})^{3}}+\frac{\sigma^{2}\lambda^{2}L^{3}(L\log(t+t_{0})+1)}{(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{\lambda^{2}L^{3}(1+L)\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(T+t_{0})^{3}}\bigg).$
@@
where $N_{t}\coloneqq\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}+\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}$. Define $G_{1}\coloneqq 240a^{2}\sigma^{2}dL$, $G_{2}\coloneqq\frac{6a^{3}\lambda^{2}L^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)}$, $G_{3}\coloneqq\frac{48a^{5}n\sigma^{2}\lambda^{2}L^{4}}{(1-\lambda)^{2}}$ and $G_{4}\coloneqq\frac{30a^{3}\lambda^{2}L^{4}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and plug into (F), to get
@@
$\epsilon_{t}^{2}=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})}{n(t+t_{0}+1)}$ $+\frac{4a^{2}\lambda^{2}L(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{32a^{4}\sigma^{2}\lambda^{2}L^{3}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{3}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}},$ (44)
@@
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)$ $=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\sigma^{2}\lambda^{2}L^{3}(L\log(t+t_{0})+1)}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{\lambda^{2}L^{3}(1+L)\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}\bigg).$

### v2

@@ 5 Conclusion
In this paper, we revisit convergence in HP of a variant of DSGD under sub-Gaussian noise. We show that DSGD is guaranteed to converge in the HP sense under the same conditions on the cost as in the MSE sense, achieving order-optimal rates for both non-convex and strongly convex costs. Moreover, our results show DSGD achieves linear speed-up in the number of users in both cases. Compared to [dsgd-high-prob], wherein the authors study HP convergence of DSGD, we relax strong conditions like uniformly bounded gradients or asymptotically vanishing noise, while showing improved rates, with linear speed-up in the number of users and transient times matching, or strictly sharper than transient times stemming from corresponding MSE rates. Compared to works studying HP convergence of centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], we provide several novel results, establishing tight control on the MGF of the consensus gap for both non-convex and strongly convex costs in Lemmas 3 and 4, a more fine-grained bound on the MGF of an “almost decreasing” process in Lemma 5, as well as showing variance reduction benefits of decentralized algorithms (in the HP sense) in Lemma 2. Future work includes extending our results to costs satisfying the PŁ condition, incorporating the GT mechanism to remove bounded heterogeneity condition and considering noise with heavier tails, like sub-Weibull.
@@ Appendix A Introduction
The Appendix contains results omitted from the main body. Section B collects some important facts used in our proofs, Section C provides some technical results, Section D defines some notions used in the analysis, Sections E and F provide proofs for non-convex and strongly convex costs, respectively, Section G provides numerical results, Section H derives the transient times, Section I provides a discussion on the novel results needed for achieving linear speed-up, Section J discusses the extension to time-varying networks, while Section K contains further discussion on dimension dependence in our bounds.
@@ Appendix B Useful Inequalities
Next, we state a useful result from [ran-improved].
Proposition 6 ([ran-improved], Lemma 21).
@@
The following result states some important consequences of conditions (A2) and (A3), see, e.g., [bertsekas-gradient, nesterov-lectures_on_cvxopt].
@@
Finally, we complete this section with the following important consequences of (A6), see, e.g., [nesterov-lectures_on_cvxopt].
@@
Let (A6) hold. Then, the following are true, for all $i\in[n]$ and $x\in\mathbb{R}^{d}$.
@@ Appendix C Technical Results
    To prove the first property, we follow steps similar to those in [li2020high, Lemma 1]. Let $y_{i}=\frac{z^{t}_{i}}{\sigma_{i}}$ and note that $\mathbb{E}\left[\exp\left(\|y_{i}\|^{2}\right)\>|\>\mathcal{F}_{t}\right]\leq\exp(1)$, from (A3). Assume first that $v\in\mathbb{R}^{d}$ is such that $\|v\|\leq\frac{4}{3}$. Using the inequality $\exp(a)\leq a+\exp(\nicefrac{{9a^{2}}}{{16}})$, which holds for any $a\in\mathbb{R}$, we then have
@@
    Combining the previous result with Lemma 1 from [jin2019short], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
@@
    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., [vershynin_2018, Proposition 2.5.2], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
@@ Appendix D Analysis Setup
We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [nedic-subgrad, jakovetic-unification, ran-improved]. Let ${\mathbf{x}}^{t}\coloneq col(x_{1}^{t},\ldots,x_{n}^{t})\in\mathbb{R}^{nd}$ denote the column vector stacking users’ local models. Using this notation, we can then represent the update rule (12) compactly as
@@
where $\mathbf{W}=W\otimes I_{d}\in\mathbb{R}^{nd\times nd}$, $\otimes$ denotes the Kronecker product and $I_{d}\in\mathbb{R}^{d\times d}$ denotes the $d$-dimensional identity matrix, while ${\mathbf{g}}^{t}\coloneqq col(g^{t}_{1},\ldots,g^{t}_{n})$. Next, define the matrix $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$, where $\mathbf{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones. The matrix $J$ represent the “ideal” consensus matrix, where all users can communicate with each other. [^13] This is equivalent to the client-server setup in terms of the update rule, as the server averages models from all clients in each iteration. The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [jakovetic-unification, ran-improved] and references therein.
@@ Appendix E Proofs for Non-convex Costs
The proof follows similar steps as in [ran-improved, Lemma 3]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
@@
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
@@
$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}$
$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
@@
    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{dLT}}$, then with probability at least $1-\delta$
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\sigma\sqrt{dL}}{\sqrt{nT}}+\frac{48n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{dLT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{dLT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma\sqrt{dL}}{\sqrt{nT}}+\frac{1}{CT}$. Therefore, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we finally have
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+120\big)}{\sqrt{nT}}$
    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{48n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
@@
    $\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{T+1}}{{\delta}})\big)}{\sqrt{n(T+1)}}$
    $+\frac{\sigma\sqrt{d}\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
    as the last term, which does not attain a linear speed-up, is no longer of higher order. We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [pmlr-v119-koloskova20a, ran-improved].
@@ Appendix F Proofs for Strongly Convex Costs
In this section we prove Lemma 4 and Theorem 2. To do so, we follow a similar strategy to the one in, e.g., [conv-rates-dsgd-jakovetic, ran-improved], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense. However, to establish guarantees in the HP sense, we instead work with the MGF of the iterates and have the following important result.
@@
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg($ $\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}$
$+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
@@
where $N_{t}\coloneqq\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}+\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}$. Define $G_{1}\coloneqq\frac{8640\kappa\sigma^{2}d}{\mu}$, $G_{2}\coloneqq\frac{1296\kappa^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+9\kappa^{2})}}{{(1-\lambda)}})}{\mu(1-\lambda)}$, $G_{4}\coloneqq\frac{6480\kappa^{3}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and $G_{3}\coloneqq\frac{373248n\kappa^{4}\sigma^{2}\lambda^{2}}{\mu(1-\lambda)^{2}}$, and plug into (F), to get
@@
$\epsilon_{t}^{2}=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})}{n(t+t_{0}+1)}$ $+\frac{144\kappa\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+9\kappa^{2})}}{{(1-\lambda)}})}{\mu(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{41472\kappa^{3}\sigma^{2}\lambda^{2}}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{720\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}},$ (44)
@@
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg($ $\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}$
$+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$